Title of Paper

The author(s)’s name(s)^{∗}

Abstract. This is to explain how to prepare a contribution for publication in EMS Congress Reports.

2010 Mathematics Subject Classification. Primary 11-XX; Secondary 14-XX. Keywords. Drinfeld modules, L-functions, Weil conjecture.

1. Introduction

Authors are requested to use standard L^{A}TEX and the class file

emsprocart.cls

This file is essentially ‘article.cls’, slightly changed, loading amsmath, amsfonts, amssymb, latexsym, and with amsthm.sty included. It sets the page size to

\textheight=548pt
\textwidth=357pt

The TEX source file should begin with

\documentclass{emsprocart}

Enter the name(s) of the author(s) using the tag

\contact[e-mail address]{Author's address}

Each author’s name should be entered with a separate \contact command. No personal style files should be used. Each paper should contain the 2000 Mathematics Subject Classification. Please DO NOT use one-letter lower case newly defined commands like

\newcommand{\e}{\varepsilon}

since this can interfere with the hyperref package and conversion of your article to Times fonts later.

2. Some rules

In order to achieve a uniform appearance of all the contributions, we encourage you to to observe the following rules when preparing your article.

∗The authors are grateful to the Max Planck Institute (Bonn) for hospitality during the writing of this paper.

2.1. Displayed formulas. If you have displayed formulas consisting of more than one line we recommend to you use

\begin{align}...\end{align}

instead of

\begin{eqnarray}...\end{eqnarray}

(respectively the starred forms) since the former yields a better spacing. Compare:

A = f(x_{i}) = F′(x)  (1)
B = g(x_{i}) = G′(x)  (2)
A = f(x_{i}) = F′(x)  (3)
B = g(x_{i}) = G′(x).  (4)

In case you do not want the numbering for every line, type

\nonumber

at the end of the line where you do not want a number.

A = f(x_{i}) = F′(x)
B = g(x_{i}) = G′(x).  (5)

If you want a number for the complete block, this works:

\begin{equation}\begin{split}...\end{split}\end{equation}

\begin{aligned}A & = f(x_{i}) = F′(x) \\ B & = g(x_{i}) = G′(x). \end{aligned}  (6)

If you prefer to number equations in the form (2.1), (2.2),…, add the line

\numberwithin{equation}{section}

to the preamble of your document. 2.2. Theorems and alike. For theorems, lemmas, definitions, etc. use the standard syntax.

\begin{theorem}...\end{theorem}

Put optional arguments into square brackets (“Main Theorem, [5]” in the example below). Theorem 2.1 (Main Theorem, [5]). If a knot K has Seifert form V_{K} and its Alexander polynomial is not 1, then there is an infinite family {K_{i}} of nonconcordant knots such that each K_{i} has Seifert form V_{K}.

Definition 2.2. A preference order (or preference relation) on X is a binary relation ≻ with the following two properties.

(1) Asymmetry: If x ≻ y, then y ⊁ x.

(2) Negative transitivity: If x ≻ y and z ∈ X, then either x ≻ z or z ≻ y or both must hold.

In this example file, enumerations of theorems, lemmas definitions, etc. appear consecutively. If you want separate numbering (Theorem 2.1, Definition 2.1) change e.g.

\newtheorem[theorem]{definition}

to

\newtheorem{definition}{Definition}[section]

If you want a statement unnumbered, just define

\newtheorem*{coro}{Corollary}

to obtain Corollary. If all the coefficients of (A.2) are entire functions, then all local solutions of (A.2) admit a meromorphic continuation over the whole complex plane ℂ.

For a proof, use

\begin{proof}...\end{proof}

An end-of-proof sign □ is set automatically. Proof. This finishes the proof of the corollary. 2.3. Operator names. There are several TEX-commands setting things automatically upright like det, sin,…. If you need operators not predefined, simply define e.g.

\newcommand{\Hom}{\operatorname{Hom}}
\newcommand{\Ker}{\operatorname{Ker}}

and then use

\Hom, \Ker

to obtain

φ ∈ Hom(G/H) ⟹ Ker(φ) ≠ {0}.

It is accepted typographical standard that abbreviated mathematical expressions standing for “words” appear in roman (upright) typeface.

3. References

It follows a list of references showing you the style in which books and journal articles should be listed.

[1] M. T. Anderson, Geometric aspects of the AdS/CFT correspondence. In AdS/CFT Correspondence: Einstein metrics and their conformal boundaries (ed. by Olivier Biquard). IRMA Lect. Math. Theor. Phys. 8, European Math. Soc. Publishing House, Zürich 2005, 1–31.

[2] I. Babuška and W. C. Rheinboldt, Error estimates for adaptive finite element computations. SIAM J. Numer. Anal. 15 (1978), 736–754.

[3] M. H. Freedman and F. Quinn, Topology of 4-manifolds. Princeton Math. Ser. 39, Princeton University Press, Princeton, NJ, 1990.

[4] A. I. Mal’cev, On homomorphisms onto finite groups. Ivanov. Gos. Ped. Inst. Ucen. Zap. 18 (1958), 49–60.

[5] M. Schweizer, Hedging of options in a general semimartingale model. Diss. ETH Zürich No. 8615, Zürich 1988.

[6] M. F. Vigneras, Induced representations of reductive p-adic groups in characteristic ℓ ≠ p. Selecta Math. (N.S) 4 (1998), 549–623.

Author’s name, Department, University, PO Box or Street, City, Country E-mail: e-mail address